OpenStudy (anonymous):

How do you create equations for this problem so I can graph it and get the answer? You bought a total of 6 pens and pencils for $4. If each pen costs $1 and each pencil costs $.50, how many pens and pencils did you buy?

OpenStudy (danjs):

was that you with this prob, no notifications to remember

OpenStudy (danjs):

no alerts of responses, sorry

OpenStudy (anonymous):

yeah lol i didn't get it

OpenStudy (danjs):

you had both lines correct 6 = x + y 4 = 1x + (1/2)y

OpenStudy (danjs):

you can just use the axis intercepts to graph them fast...

OpenStudy (danjs):

6 = x + y if x=0, y = 6 if y = 0, x = 6, 2 points, there is the line

OpenStudy (danjs):

same for the other line

OpenStudy (anonymous):

hmm, but i got 2 pens and 4 pencils for the answer

OpenStudy (danjs):

yes, you set both lines equal to solve that, on the graph, both lines are equal where they intersect one another at a point (x,y)

OpenStudy (danjs):



OpenStudy (danjs):

the scales are messed up, but solution should be at (2,4)

OpenStudy (anonymous):

ohh wow, thanks so much

OpenStudy (danjs):

i just limited it to the first quadrant, since pens and pencils are positive numbers

OpenStudy (anonymous):

i see , now i get it thanks!

OpenStudy (danjs):

welcome